Use the compound interest formula A=P(1+r/n)^nt to find approximately how many years an investment of $14,000 will accumulate to $42,000 if it earns 8% compounded quarterly. thanks.
Answer by stanbon(75887)

Use the compound interest formula A=P(1+r/n)^nt to find approximately how many years an investment of $14,000 will accumulate to $42,000 if it earns 8% compounded quarterly. thanks
-------------------
A(t) = P(1+ (r/n))^(nt)
42000 = 14000(1 + (0.08/4))^(4t)
3 = 1.02^(4t)
4tlog(1.02) = log3
4t = 55.478...
t = 13.87 years
========================
Cheers,
Stan H.